OpenStudy (hoybaby):

how do i simplyify or this equation? (3m6g15)^3

OpenStudy (hoybaby):

(3m^6g1^5)^\[(3m ^{6}g ^{15})^{3}\]3

OpenStudy (saifoo.khan):

\[\Large (3)^3 \times m^{6\times 3} \times g^{15 \times 3}\]

OpenStudy (saifoo.khan):

\[\Large 27m^{18}g^{45}\]

OpenStudy (hoybaby):

can you help me with another one? i do not get the hang of this

OpenStudy (saifoo.khan):

umm, okay. lol.

OpenStudy (hoybaby):

\[5x ^{2}\sqrt{3x ^{12}} \times 9\sqrt{3x ^{20}}\]

OpenStudy (hoybaby):

?

OpenStudy (saifoo.khan):

u here?

OpenStudy (hoybaby):

yea

OpenStudy (hoybaby):

i sent the equation

OpenStudy (saifoo.khan):

solving..

OpenStudy (hoybaby):

i am totally confused on simplying these radicals like this

OpenStudy (saifoo.khan):

Do u want the simplified answer? if u are doing online thing

OpenStudy (hoybaby):

\[\sqrt{16}x ^{6} \over \sqrt{448}\] this stumbles me. the equation i sent well i tried it and i got it wrong on my webstie called studyisland

OpenStudy (saifoo.khan):

\[\Large 135x^{18}\] this is for the one about this..

OpenStudy (saifoo.khan):

\[\sqrt{16x ^{6}} \over \sqrt{448}\]is that like this?

OpenStudy (hoybaby):

the 135x^18 is the answer for the other one but what is the answer and how do you get it for the one you sent \[\sqrt{16x ^{2}} \over \sqrt{448}\]

OpenStudy (hoybaby):

i dont know how to solve problems like this, can you help me with the processes so i can do it by myself?

OpenStudy (saifoo.khan):

\[5x ^{2}\sqrt{3x ^{12}} \times 9\sqrt{3x ^{20}}\] \[(5*9)x^2 \sqrt{9x^{32}}\] \[45x^2*3 \sqrt{ x^{32}}\] \[135x^{2+16}\] \[135x^{18}\]

OpenStudy (hoybaby):

how about \[\sqrt{16x ^{2}}\over \sqrt{448}\]

OpenStudy (saifoo.khan):

i wil be back in 5mins.. can u wait please?

OpenStudy (hoybaby):

i can wait

OpenStudy (hoybaby):

i mean, this is proper problem \[\sqrt{16x ^{6}}\over \sqrt{448}\]

OpenStudy (hoybaby):

?

OpenStudy (saifoo.khan):

\[\sqrt{\frac{16x^6}{448}}\] \[\sqrt{\frac{1x^6}{28}}\] \[\frac{x^3}{2\sqrt7}\]answer

OpenStudy (hoybaby):

how did you get that?

OpenStudy (saifoo.khan):

i showed u the working..

OpenStudy (hoybaby):

how did you get 1x6^

OpenStudy (hoybaby):

i mean 1x^6

OpenStudy (saifoo.khan):